# Gaia Data Release 3: spectroscopic binary-star orbital solutions

### v1

@@ Acknowledgements.
- Pourbaix et al. (2022) Pourbaix, D., Arenou, F., Gavras, P., et al. 2022, Gaia DR3 documentation Chapter 7: Non-single stars, Gaia DR3 documentation, European Space Agency; Gaia Data Processing and Analysis Consortium. Online at ¡A href=“https://gea.esac.esa.int/archive/documentation/GDR3/index.html”¿ https://gea.esac.esa.int/archive/documentation/GDR3/index.html¡/A¿, id. 7
@@
- Price-Whelan et al. (2020a) Price-Whelan, A. M., Hogg, D. W., Rix, H.-W., et al. 2020a, ApJ, 895, 2
- Price-Whelan et al. (2020b) Price-Whelan, A. M., Hogg, D. W., Rix, H.-W., et al. 2020b, ApJ, 895, 2
@@
- Sartoretti et al. (2022) Sartoretti, P., Blomme, R., David, M., & Seabroke, G. 2022, Gaia DR3 documentation Chapter 6: Spectroscopy, Gaia DR3 documentation, European Space Agency; Gaia Data Processing and Analysis Consortium. Online at ¡A href=“https://gea.esac.esa.int/archive/documentation/GDR3/index.html”¿ https://gea.esac.esa.int/archive/documentation/GDR3/index.html¡/A¿, id. 6
@@ Appendix A Historical developments
Once the period evaluated, the general model to fit (Eq. 1) is still non-linear in some of the parameters (mainly the eccentricity ). This implies that the of the fit could present a few or even numerous minima in the parameter space, and to find the global one is a true challenge. In particular, the derivation of the eccentricity is a notorious pitfall. The search for the global minimum might be performed empirically by guessing its approximate position and then by descending the valley by classical numerical methods (such as the Levenberg-Marquardt algorithm). However, in a pipeline as the one developed for Gaia, identifying the location of the global minimum is uncertain and a specific type of search is necessary. This is particularly frustrating because a well-trained expert eye is able to estimate a rough value for the eccentricity with great efficiency. In this context, we explored the possibilities to adopt alternative approaches. We acted along two main axes. On one hand, we tried to invent a method based on pattern recognition in the phase diagram of the RVs folded with the candidate periods. Such an approach would allow having a first guess for the eccentricity, which could be used for further classical derivation. In order to apply pattern recognition techniques, it is mandatory to define an interpolation function in the phase diagram in order to obtain a robust description of the shape of the RV curve. Several tools have been explored such that the nearest-neighbour interpolation, the linear interpolation, the interpolation by cosines, the local Akima periodic interpolation, the periodic interpolation by cubic splines, the periodic interpolation by Catmull-Rom splines. We also explored the possibilities of performing classical adjustment by least-square methods. In particular, we considered classical polynomials, Fourier series, Bézier curves (using De Casteljan algorithm) and finally, cubic B-splines in their open uniform version. Once the interpolation curve is chosen and the shape is defined, we tested various pattern recognition methods such as algorithms for regression based on the nearest neighbourhood, algorithms for regression based on trees and in particular on extremely randomised trees and, finally, algorithms for regression based on neural networks. Full details of this exploration and detailed references are given in the work of Delchambre (2010).
@@
On the other hand, we developed an approach based on genetic algorithms. These algorithms have the advantage to circumvent any preliminary guess of the solution. This is at the expense of being obliged to explore (even cleverly) the space of the parameters using the generation of random numbers (Monte Carlo methods). This work is described in details in Delchambre (2010). The use of genetic algorithms is a trade-off between the probability to effectively detect the global minimum and the consumption of computer time. The code using genetic algorithms were written, delivered and tested. The detection of the global minimum was possible but was implying a prohibitively large amount of computer time despite strong efforts made to accelerate it. The large amount of objects to treat in the Gaia processing suggests using much more efficient algorithms. The genetic algorithm code was not activated for the operational runs.
@@
Other methods are now available to derive orbital solutions. They became widespread in the domain of exoplanets, where the orbits are particularly incompletely covered and complex. The majority of these methods are based on Monte Carlo simulations. Very often, they use the Markov Chain Monte Carlo (MCMC) method (see e.g. Sharma 2017) often utilised in conjunction with Bayesian approaches. These algorithms necessitate a dense coverage of the full space of priors, at the expense of the computer time. Comparing the works of Price-Whelan et al. (2017) and Price-Whelan et al. (2020b) pinpoints the strong impact that could have the choice of priors on the resulting posteriors. The Bayesian methods have admittedly the big advantage to potentially provide marginal posterior probabilities for a series of models and thus estimate how much the observational data could favour a particular model. According to Ford & Gregory (2007), the challenge of estimating marginal posterior probabilities is still a crucial problem and obvious estimators are suffering from poor convergence properties. In addition, for the preliminary DR3 catalogue, the directives were that we should output one and only one model per treated object, thus mitigating the interest of using Bayesian methods. In addition, these Monte Carlo methods are computer time-consuming and were not acceptable for the DR3 processing.
We finally adopted a method based on fully analytical and classical numerical developments.
@@ Appendix B Practical details on the processing
The Data Processing Centre at CNES (DPCC181818https://www.cosmos.esa.int/web/gaia/dpac) is located at the Centre National d’Etudes Spatiales in Toulouse (France). It ran the spectroscopic processing chain (CU6) and the object processing chain (CU4). Globally, it ran the whole set of NSS codes and, in particular, the spectroscopic orbital solution chain described in the present paper. Our NSS-SB1 processing chain delivers the results to the combiner. The Main Data Base (MDB) is fed with solutions from this combiner. If the combiner finds no other solutions to combine with, the spectroscopic results are directly transmitted to the MDB. If a combination is possible, only the output of the combination is saved.
@@
Appendix C The separable variable technique applied to the Keplerian model fit
@@ C.1 A reminder of the Levenberg-Marquardt method
being the array of 6 parameters where the superscript denotes the transpose operation. The Newton-Raphson algorithm is based on the Taylor first order development
@@
The objective function to be minimised is the quadratic sum of residuals
@@
leading in the case of Newton-Raphson to the algebraic system
@@
is the damping factor, chosen to be large when approaches the singularity.
C.2 The separable variable technique
The parameters are linear because they satisfy the property
Hereafter, subscripts and refer to linear and non-linear respectively. is the Jacobian matrix whose columns are the derivatives of with respect to .
While in both Newton-Raphson and the (commonly used) Levenberg-Marquardt (LM) algorithms, linear and non-linear parameters are treated the same way, the separable variable technique has the advantage that can be obtained for every non-linear parameter by a linear least-square regression, such that 191919The inversion of the normal equations is known to be numerically unstable and should be avoided in favour of orthogonal decomposition approaches.
@@ C.2 The separable variable technique
Therefore, Eq. 22 becomes a function of only
@@
Using Eq. 31 and Eq. 32, we deduce the derivatives
@@
with being the new hessian matrix of the system. We have then split the 66 system (Eq. 28) into two smaller systems, each one of 33, in Eq. 30 and Eq. 39 respectively.
@@ Appendix G Processing steps and outcome of analysis for SB9 systems
For more than half of the remaining systems, the CU6 data are either not transmitted to our pipeline for analysis or will not lead to any solutions. As discussed in Sects. 2 and 6.3.1, it is mainly because there is either no compelling evidence for variability in the RV time-series, there are not enough valid RV measurements, or the source lies outside the or ranges. We note that 20% of the entries in the whole SB9 contain an OBA component. -
@@
Appendix H Parameters delivered to the catalogue for classes SB1, SB1C and TrendSB1.
@@
| Transfer from CU6 products | RV time-series available | Katz et al. (2023) |
@@
| objects/time-series at entry | in [3875 - 8125] K | Sect. 2 |
@@
| RV outlier rejected from time-series | Sect. 6.3.1 | |

### v2

@@ Acknowledgements.
- Pourbaix et al. (2022) Pourbaix, D., Arenou, F., Gavras, P., et al. 2022, Gaia DR3 documentation Chapter 7: Non-single stars, Gaia DR3 documentation, European Space Agency; Gaia Data Processing and Analysis Consortium. Online at https://gea.esac.esa.int/archive/documentation/GDR3/index.html, 7
@@
- Price-Whelan et al. (2020) Price-Whelan, A. M., Hogg, D. W., Rix, H.-W., et al. 2020, ApJ, 895, 2
@@
- Sartoretti et al. (2022) Sartoretti, P., Blomme, R., David, M., & Seabroke, G. 2022, Gaia DR3 documentation Chapter 6: Spectroscopy, Gaia DR3 documentation, European Space Agency; Gaia Data Processing and Analysis Consortium. Online at https://gea.esac.esa.int/archive/documentation/GDR3/index.html, 6
@@ Appendix A Historical developments
Once the period has been evaluated, the general model to fit (Eq. 1) remains non-linear in some of the parameters (mainly the eccentricity, ). This implies that the of the fit could present a few or even numerous minima in the parameter space, and to find the global one is a true challenge. In particular, the derivation of the eccentricity is a notorious pitfall. The search for the global minimum might be performed empirically by guessing its approximate position and then by descending the valley by classical numerical methods (such as the Levenberg-Marquardt algorithm). However, in a pipeline as the one developed for Gaia, identifying the location of the global minimum is uncertain and a specific type of search is necessary. This is particularly frustrating because a well-trained expert eye is able to estimate a rough value for the eccentricity with great efficiency. In this context, we explored the possibilities to adopt alternative approaches. We acted along two main axes. On the one hand, we tried to invent a method based on pattern recognition in the phase diagram of the RVs folded with the candidate periods. Such an approach would allow us to take a first guess for the eccentricity, which could then be used for further classical derivation. In order to apply pattern recognition techniques, it is mandatory to define an interpolation function in the phase diagram to obtain a robust description of the shape of the RV curve. Several tools have been explored such that the nearest-neighbour interpolation, the linear interpolation, the interpolation by cosines, the local Akima periodic interpolation, the periodic interpolation by cubic splines, and the periodic interpolation by Catmull-Rom splines. We also explored the possibilities of performing classical adjustment by least-square methods. In particular, we considered classical polynomials, Fourier series, Bézier curves (using De Casteljan algorithm) and finally, cubic B-splines in their open uniform version. Once the interpolation curve is chosen and the shape is defined, we tested various pattern recognition methods such as algorithms for regression based on the nearest neighbourhood, algorithms for regression based on trees and in particular on extremely randomised trees and, finally, algorithms for regression based on neural networks. Full details of this exploration and detailed references are given in the work of Delchambre (2010).
@@
On the other hand, we developed an approach based on genetic algorithms. These algorithms have the advantage to circumvent any preliminary guess of the solution. This is at the expense of being obliged to explore (even cleverly) the space of the parameters using the generation of random numbers (Monte Carlo methods). This work is described in details in Delchambre (2010). The use of genetic algorithms is a trade-off between the probability to effectively detect the global minimum and the consumption of computer time. The codes using genetic algorithms were written, delivered and tested. The detection of the global minimum was possible but was implying a prohibitively large amount of computer time despite strong efforts made to accelerate it. The large amount of objects to treat in the Gaia processing suggests using much more efficient algorithms. The genetic algorithm code was not activated for the operational runs.
@@
Other methods are now available to derive orbital solutions. They became widespread in the domain of exoplanets, where the orbits are particularly incompletely covered and complex. The majority of these methods are based on Monte Carlo simulations. Very often, they use the Markov chain Monte Carlo (MCMC) method (see e.g. Sharma 2017) often utilised in conjunction with Bayesian approaches. These algorithms necessitate a dense coverage of the full space of priors, at the expense of the computer time. Comparing the works of Price-Whelan et al. (2017) and Price-Whelan et al. (2020) pinpoints the strong impact that could have the choice of priors on the resulting posteriors. The Bayesian methods have admittedly the big advantage to potentially provide marginal posterior probabilities for a series of models and thus estimate how much the observational data could favour a particular model. According to Ford & Gregory (2007), the challenge of estimating marginal posterior probabilities is still a crucial problem and obvious estimators are suffering from poor convergence properties. In addition, for the preliminary DR3 catalogue, the directives were that we should output one and only one model per treated object, thus mitigating the interest of using Bayesian methods. In addition, these Monte Carlo methods are computer time-consuming and were not acceptable for the DR3 processing. Finally, we adopted a method based on fully analytical and classical numerical developments.
@@ Appendix B Practical details on the processing
The Data Processing Centre at CNES (DPCC191919https://www.cosmos.esa.int/web/gaia/dpac), located at the Centre National d’Etudes Spatiales in Toulouse (France), ran the spectroscopic processing chain (CU6) and the object processing chain (CU4). Globally, it ran the whole set of NSS codes and, in particular, the spectroscopic orbital solution chain described in the present paper. Our NSS-SB1 processing chain delivers the results to the combiner. The Main Data Base (MDB) is fed with solutions from this combiner. If the combiner finds no other solutions to combine with, the spectroscopic results are directly transmitted to the MDB. If a combination is possible, only the output of the combination is saved.
@@
Appendix C Separable variable technique applied to the Keplerian model fit
@@ C.1 A reminder of the Levenberg-Marquardt method
with being the array of 6 parameters where the superscript denotes the transpose operation. The Newton-Raphson algorithm is based on the Taylor first order development
@@
The objective function to be minimised is the quadratic sum of residuals as
@@
In the case of Newton-Raphson, this leads to the following algebraic system:
@@
where is the damping factor, chosen to be large when approaches the singularity.
C.2 Separable variable technique
The parameters are linear because they satisfy the following property:
Hereafter, subscripts and refer to linear and non-linear respectively. Then, is the Jacobian matrix whose columns are the derivatives of with respect to .
While in both Newton-Raphson and the (commonly used) Levenberg-Marquardt (LM) algorithms, the linear and non-linear parameters are treated the same way, the separable variable technique has the advantage that can be obtained for every non-linear parameter by a linear least-square regression, such that 202020The inversion of the normal equations is known to be numerically unstable and should be avoided in favour of orthogonal decomposition approaches.
@@ C.2 Separable variable technique
Therefore, Eq. 22 becomes a function of only, such that
@@
Using Eq. 31 and Eq. 32, we deduce the derivatives as
@@
with being the new hessian matrix of the system. We have then split the 66 system (Eq. 28) into two smaller systems, each one of 33, in Eq. 30 and Eq. 39, respectively.
@@ Appendix G Processing steps and outcome of analysis for SB9 systems
For more than half of the remaining systems, the CU6 data are either not transmitted to our pipeline for analysis or will not lead to any solutions. As discussed in Sects. 2 and 6.3.1, it is mainly because there is either no compelling evidence for variability in the RV time series, there are not enough valid RV measurements, or the source lies outside the or ranges. We note that 20% of the entries in the whole SB9 contain an OBA component. -
@@
Appendix H Parameters delivered to the catalogue for classes SB1, SB1C and TrendSB1
@@
| Transfer from CU6 products | RV time series available | Katz et al. (2023) |
@@
| objects/time series at entry | in [3875 - 8125] K | Sect. 2 |
@@
| RV outlier rejected from time series | Sect. 6.3.1 | |